OpenStudy (anonymous):

is -7+9=-9+7 true, false, or open? plz help A: True B: False C: Open

OpenStudy (anonymous):

Well lets see if the right hand side of the equation equals the left hand side of the equal Left Side: -7+9 = 2 Right Side: -9+7=-2 Does 2=-2???//?

OpenStudy (kamibug):

It would be true if the expression on the right was 9-7, instead. :P

OpenStudy (anonymous):

Looks false. :) 2 does not = to -2

OpenStudy (kamibug):

\(\huge\cal\color{royalblue}{Welcome~to~Openstudy!~:)}\) @Katy_kat18